Use Theorems about sines and cosines to prove that sin(pi-theta) = cos((pi/2)-theta)
Answer by stanbon(75887)

Use Theorems about sines and cosines to prove that sin(pi-theta) = cos((pi/2)-theta)
-------
sin(pi)*cos(t)-cos(pi)sin(t) = sin(t)
------
0*cos(t) - (-1)sin(t) = sin(t)
sin(t) = sin(t)
-----------------------
Cheers,
Stan H.
=======================